OpenStudy (anonymous):

how do you do elimination using multiplication ? example problem : 3x-3y=-6 -5x+6y=12

OpenStudy (anonymous):

Multiply the first by 5 and the second by 3.

OpenStudy (anonymous):

multiply the first equation with 2 ,then add the two equations y will be eliminated

OpenStudy (anonymous):

Well, in order to do elimination, first you have to change one of the equations using multiplication. 3x-3y=-6 2(3x-3y) = 2(-6) 6x-6y= -12 Now, elimination: 6x-6y= -12 + -5x+6y=12 --------------------- ?

OpenStudy (anonymous):

3x-3y=-6 (*5), and -5x+6y=12 by (*3)